determine the ratio of substance needed to prepare a carbonate buffer of pH= 10.50 providing the available substances are H2CO3, NaHCO3, and Na2CO3. given these equilibrium systems: use Henderson-Hasselbach equation for pH.

H2CO3 <------> H^+ +HCO^3-
Ka1 = 4.2x10^-7 [H^+] =???
Answer #1

[ Na2CO3] = concentration of Na2CO3

[ HCO3-] = cocncentration of NaHCO3 used

Given , Ka1 = 4.2 x 10-7

                 pH = 10.50

             pKa1 = - log Ka1 = - log (4.2 x 10-7 ) = 6.38

             As pKa1 is very far from needed pH so we consider pKa2

              Ka2 = 4.69 x 10-11

              pKa2 = -log ( Ka2 ) = - log ( 4.69 x 10-11 ) = 10.33

As pKa2 is near required pH we use a buffer of sodium carbonate and sodium hydrogen carbonate

Henderson-Hasselbach equation ,

Molar Ratio of sodium carbonate to sodium bicarbonate = 1.48 : 1

Weight ratio of Na2CO3 to NaHCO3 = 1.48 x [ (molar weight of Na2CO3 ) / ( molar weight of NaHCO3 ) ]

                                                          = 1.48 x ( 106 / 84 )

                                                         = 1.87 : 1